1. Build up the operation tables for group G with orders 1, 2, 3 and 4 using the elements a, b, c, and e as the identity element in an appropriate way.
2. i. State the Lagrange’s theorem of group theory.
ii. For a subgroup H of a group G, prove the Lagrange’s theorem.
iii. Discuss whether a group H with order 6 can be a subgroup of a group with order 13 or not. Clearly state the reasons.
1
Expert's answer
2020-04-27T19:42:09-0400

1.For the order 1 we obtain, that |G| = 1. Therefore G = {e}.

For the order 2 we obtain, that |G| = 2. Therefore G= {e,a}, where a2 = 1. The respective table is:

"e \cdot e = e"

"e\cdot a = a"

"a \cdot e = a"

"a \cdot a = e"

For the order 3 we obtain, that |G| = 3. Therefore G={e,a,b}, where a3 = b3 = e and a-1 = b. The respective table is:

"e \cdot e = e" ; "a \cdot e = a" ; "b \cdot e = b"

"e \cdot a = a" ; "a \cdot a = a^2 = a^3 \cdot a^{-1} = e \cdot b = b" ; "b\cdot a = a^{-1} \cdot a = e"

"e\cdot b = b" ; "a \cdot b = a \cdot a^{-1} = e" ; "b \cdot b = b^2 = a^{-2} = a^{-2} \cdot a^3 = a"

For the order 4 we obtain, that |G| = 4. Therefore G = {e,a,b,c}. Since there are 2 non-isomorphic groups of the order 4, we will consider the cyclic one. This means, that:

"e \cdot e = e" ; "a \cdot e = a" ; "b \cdot e = b" ; "c \cdot e = c"

"e \cdot a = a" ; "a \cdot a = a^2 = b" ; "b \cdot a = a^2 \cdot a = a^3 = c" ; "c \cdot a = a^3 \cdot a = e"

"e \cdot b = b" ; "a \cdot b = a \cdot a^2 = a^3 = c" ; "b \cdot b = b^2 = a^4 = e" ; "c \cdot b = a^3 \cdot a^2 = a"

"e \cdot c = c" ; "a \cdot c = a \cdot a^3 = a^4 = e" ; "b \cdot c = a^2 \cdot a^3 = a^5 = a^1 = a" ; "c \cdot c = a^3 \cdot a^3 = a^2 = b"

2.1)Let's state the Lagrange's theorem:

Let H be a subgroup of group G. Then |G| = [G:H] "\cdot" |H|, where [G:H] is the number of right cosets of H in G.

2) For the proof we will consider the right cosets of H in G. Indeed, for any aH and bH either aH"\cap" bH = "\emptyset" or aH = bH. this statement is true due to the right cosets are the equivalent classes, and since if a,b belong to the same class then a-1b belong ot H. So, if c belong to aH and bH, then a-1c and c-1b belong to H. Thus "a^{-1} \cdot c \cdot c^{-1} \cdot b = a^{-1} \cdot b \in H" . So aH = bH. Now, since |aH| = |H| for every a from G(since f : H -> aH such that f(x) = ax is clearly bijective), we obtain that |G| = [G:H] "\cdot" |H|, where [G:H] is the number of right cosets of H in G.

3) It is not true. As an obvious consequence of this theorem, we obtain that if |H| = 6, |G| "\neq" 13, since 6 doesn't divide 13.